# ViG-Bias: Visually Grounded Bias Discovery and Mitigation

### v1

@@ Abstract
The proliferation of machine learning models in critical decision making processes has underscored the need for bias discovery and mitigation strategies. Identifying the reasons behind a biased system is not straightforward, since in many occasions they are associated with hidden spurious correlations which are not easy to spot. Standard approaches rely on bias audits performed by analyzing model performance in pre-defined subgroups of data samples, usually characterized by common attributes like gender or ethnicity when it comes to people, or other specific attributes defining semantically coherent groups of images. However, it is not always possible to know a-priori the specific attributes defining the failure modes of visual recognition systems. Recent approaches propose to discover these groups by leveraging large vision language models, which enable the extraction of cross-modal embeddings and the generation of textual descriptions to characterize the subgroups where a certain model is underperforming. In this work, we argue that incorporating visual explanations (e.g. heatmaps generated via GradCAM or other approaches) can boost the performance of such bias discovery and mitigation frameworks. To this end, we introduce Visually Grounded Bias Discovery and Mitigation (ViG-Bias), a simple yet effective technique which can be integrated to a variety of existing frameworks to improve both, discovery and mitigation performance. Our comprehensive evaluation shows that incorporating visual explanations enhances existing techniques like DOMINO, FACTS and Bias-to-Text, across several challenging datasets, including CelebA, Waterbirds, and NICO++.
@@ 1 Introduction
The increasing adoption of deep learning-based image classification models, notably in systems like facial recognition software, raises concerns related to biases and fairness[3]. Biases in such systems can lead to unfair and discriminatory outcomes, making it crucial to identify and address their underlying causes. Unfortunately, determining on which subgroups a given system may exhibit biased behavior is not an easy task. The typical approach to detect biases consist in auditing the system with respect to pre-defined attributes (often referred as protected attributes), like gender, sex, age or ethnicity, by comparing specific fairness metrics [2]. However, attributes like sex, age, or ethnicity are not always the real causes that explain the poor performance of a given model in a particular set of data samples. For example, in face recognition, a system may systematically fail to classify people with glasses due to the fact that they were not well represented in the training set. In this case, the auditing with standard demographic attributes, may not be able to unveil the real reason of bias. In other cases, such as the classification of animals or objects, the attributes that characterize the under-performing groups may not be obvious. Spurious correlations could lead to systems that make incorrect predictions because the model has learned these false relationships from the training data [8], or predictions which are actually correct but for the wrong reasons [24]. This issue often arises when the model is trained on datasets that do not accurately represent the real-world scenarios on which it is intended to work with, or when the dataset itself contains hidden biases. Thus, discovering the failure modes and providing explanations useful to understand the real reasons behind them becomes a crucial task when designing robust and fair systems. Existing literature refers to this task as unsupervised discovery of bias [14], or slice discovery [7], in the sense that such methods aim at mining the input data for semantically meaningful subgroups (slices) on which the model performs poorly. In this context, a slice is defined as a group of data samples that share a common attribute or characteristic which is not related to the target label.
@@
Recent studies have shown that text explanations [12] produced by multimodal large vision language models (VLMs) [6] can assist in discovering and describing subpopulations on which a model systematically underperforms, without the need for protected attribute annotations. At the same time, a different line of work [14] has qualitatively demonstrated that visual explanations [1] derived from biased models, especially for instances originating from conflicting data slices, predominantly highlight spurious features usually associated with shortcut learning. In this work, we hypothesize that, since visual explanations for biased models tend to underscore spurious correlations, they can serve as instrumental tools in informing VLMs when they are used to identify and mitigate undesired biases. To this end, we leverage the descriptive power of recent multimodal VLMs, and combine it with the capabilities of visual explanation mechanisms to uncover biases contributing to systematic failures for unknown slices in visual recognition models. Our extensive empirical evaluation highlights the advantages of the proposed method, being easily adapted to different bias discovery frameworks and boosting systematically their performance.
@@ 2 Related work
Visual explanations. Visual explanations in computer vision have largely contributed to the interpretability and trustworthiness of deep learning models. They provide insights into the decision-making process of neural networks by highlighting regions within an input image that are significant for predictions. One prominent technique in this domain is Gradient-weighted Class Activation Mapping (Grad-CAM) [21]. Grad-CAM utilizes the gradients flowing into the final convolutional layer of a CNN to understand which features are most important for a particular decision. By creating a heatmap of these weighted features, Grad-CAM offers a visual representation that demonstrates how the model focuses on different parts of the image to make a decision. This method not only aids in improving model transparency but also helps in diagnosing potential reasons for misclassifications. GradCAM has previously been employed as a description mechanism to interpret failure modes in [14]. However, even though the work of Krishnakumar et al [14] shows the potential of visual explanations for this task, they mostly provide qualitative examples and do not propose an automatic approach to bias discovery and mitigation. Singla and co-workers [22] present another interesting work where they generate annotations for visual explanations of the ImageNet dataset. They released a dataset called Salient ImageNet, which provides heatmaps distinguishing between spurious and core features, where the last ones correspond to a set of visual features that are always a part of the object definition. Even though they employ human supervision to identify the visual explanations related to spurious correlations (making the process difficult to automate), this work shows that such heatmaps can actually be used to pinpoint this type of spurious attributes. Here, we plan to leverage GradCAM heatmaps as a mechanism to direct the attention of bias discovery and mitigation methods towards spurious correlations.
@@
Automatic bias discovery using VLMs. Several recent efforts have been conducted in the identification of biased data slices using cross-modal embeddings. The DOMINO method [7] combines cross-modal embeddings and error-aware mixture models to find semantically coherent clusters of images where a given classification system is failing. Textual descriptions derived from the cross-modal embeddings are then used to characterize the failure modes that were just discovered. A similar approach is introduced by the FACTS (First Amplify and then Slice to Discover Bias) [27] method, where the authors propose to first amplify the sources of bias in the training to ease its identification. This process is carried out in two stages. The first stage involves amplifying the bias by regularizing the model. This is achieved through regularization by specifically increasing the weight decay factor $\lambda$, which penalizes the more significant weights in the model. Regularization is a common technique used to prevent overfitting by discouraging complex models, but in this context, it serves a dual purpose: it not only helps to prevent overfitting but also forces the model to rely more on the spurious attribute by simplifying the hypothesis space. The second stage consists of identifying underperforming data segments that exhibit unique correlations by employing mixture modeling within a feature space aligned with bias, a method known as correlation-aware slicing, similar to the aforementioned DOMINO. A different approach titled Bias-to-Text (B2T) was introduced in [12]. B2T uses linguistic interpretation to identify and mitigate biases in vision models, including image classifiers. By generating linguistic descriptions of images using VLMs, B2T extracts keywords indicative of bias from the wrongly predicted examples, thereby enabling a more focused understanding and correction of prejudices within the models.
@@
Bias mitigation: improving robustness to spurious correlations. Discovering conflicting slices is important, but mitigating such biases is also crucial. A common approach is based on group distributionally robust optimization (GroupDRO) [20]. GroupDRO aims to enhance model performance across different predefined groups or subpopulations in the data, based on pre-defined protected attributes like race or gender. The objective is to minimize the worst-case loss across these groups, thereby encouraging fairness and reducing disparities in model performance among them. However, one notable drawback of this technique is the requirement to assign specific group labels to every dataset entry, which can lead to substantial increases in annotation costs, and requires to know before-hand such groups. Another existing approach in the bias mitigation field is JTT (Just Train Twice) [15]. JTT starts with the standard training of the model to pinpoint examples that have been misclassified. Following this, the training process is adjusted by reweighting the dataset, giving greater importance to these previously misclassified examples. JTT forces the model to pay more attention to examples where the model is failing, thereby enhancing its performance on those specific examples. Here, we will employ JTT as a baseline approach for comparison, and propose ways to integrate visual explanations into the group definitions for GroupDRO.
@@ 3 Preliminaries and Problem Formulation
The dataset $\mathcal{D}$ is partitioned into different and not equally distributed groups or slices. A group is defined based on the combination of the target labels $y\in\mathcal{Y}$ and the spurious attributes $a_{i}\in\mathcal{A}$ that spuriously correlates with the label (i.e. $\mathcal{G}=\mathcal{A}\times\mathcal{Y}$). Formally, given a pair $(a_{i},y^{\star})$, we define a group or slice $g(a_{i},y^{\star})$ as the set of data samples labeled with $y^{\star}$ which contain attribute $a_{i}$, i.e. where $a_{i}(x)=1$.
For example, in the CelebA dataset[17], where the task involves classifying people with blond hair, the spurious attribute is the person’s gender, as there is a higher frequency of blond women in the dataset. Figure 1 illustrates this issue for the CelebA dataset, and a similar case found in the Waterbirds dataset [20], highlighting the classes with the least training samples per group.
@@ 4 ViG-Bias: Visually Grounded Bias Discovery and Mitigation
Consider a dataset $D$ with samples $(x,y,a)\in\mathcal{D}$ drawn from $\mathcal{X}\times\mathcal{Y}\times\mathcal{A}$, where $a$ is spuriously correlated with the label $y$. We are interested in discovering the visual features upon which a model have relied to make a prediction. Prior work [11, 22] suggests that, under the presence of spurious correlations, visual explanations for models trained with standard empirical risk minimization (ERM) tend to focus on such shortcuts, while ignoring the core features. We thus resort to GradCAM heatmaps generated using an image classifier that was trained with standard ERM on the dataset of interest $\mathcal{D}$. Such heatmaps will then be used as visual explanations to guide the attention of bias discovery methods (alternative visual explanation methods like ScoreCAM [26], FullGrad [23], GradCAM++ [4] are also evaluated in Section 6.3). GradCAM heatmaps are obtained by first computing the gradients of the target class with respect to the feature maps of the last convolutional layer, as neurons in this layer offer best compromise between high-level semantics and detailed spatial information [21]. These gradients flowing back are then globally-averaged over the width and height dimensions to obtain the importance weights of every feature map in this last layer. Then, the ReLU activation function is applied on the weighted average of the feature maps, so that only the features that have a positive influence on the class of interest are finally highlighted. These heatmaps are finally normalized between 0 and 1.
@@ 4.1 Are Visual Explanations Focusing on Spurious Correlations?
The work of Krishnakumar [14] and colleagues provides qualitative examples showing that GradCAM heatmaps for biased models tend to focus on spurious correlations. Here we perform a more systematic quantitative evaluation to confirm this hyphotesis. To this end, we design a simple experiment using the CelebA (where the conflicting slices are well know) that consists of three steps:
@@
    We first construct binary segmentation masks $m^{c}_{i}$ and $m^{s}_{i}$ for all images $x_{i}$ in both datasets, segmenting the core features associated to the object of interests ($m^{c}_{i}$ representing hair in CelebA), and the spurious attribute ($m^{s}_{i}$ representing the face) as shown in Figure 1. We do so by using LangSAM  [^1] https://github.com/luca-medeiros/lang-segment-anything, an open-source model that combines Segment Anything Model (SAM) [13] with GroundingDINO [16] to enable instance segmentation from text prompts.
@@
Figure 2: Visual explanation methods (e.g. GradCAM[21] for the class blond) help with identifying spurious correlations. We use the Intersection-over-Union (IoU) metric to measure the percentage of overlap between a binary image representing the spurious feature and the image we get after applying a visual explanation methods.
@@ 4.2 Improving Unsupervised Bias Discovery via Visual Explanations
We modify DOMINO by simply pre-processing the images before the cross-modal embeddings are obtained, i.e. we generate heatmaps using the pretrained classifier and extract embeddings for $h(x)$ instead of $x$. Note that other type of visual explanation methods could also be used, as discussed in Section 6.3.
Visually Grounded FACTS (ViG-FACTS): The FACTS method is similar to DOMINO, but it consists of a two-stage process, where biases are first amplified and then slicing is performed. In the first stage, we increase the model’s reliance on spurious, often bias-indicative correlations, which are amplified using regularization. In general, regularization helps in preventing overfitting and encourage the model to learn more robust patterns that are assumed to be applicable beyond the training set. However, in FACTS, the regularization is applied to constrain the model’s capacity and learn the bias-aligned slices where spurious correlation holds. This leads to the model developing a strong dependency on these correlations, making the biases more pronounced and identifiable. Once these biases are amplified and more clearly delineated, the second stage of the process, slicing, an error-aware mixture model that uses the CLIP embeddings similarly to DOMINO is applied.
@@
Figure 4: Our objective is to pinpoint slices of data where a spurious correlation exists between a task-irrelevant attribute (such as gender) and the task label of interest. For instance, in the given example, above, this corresponds to not blond in male or female, while not blond male forms a bias-aligned slice. FACTS amplifies correlations, aiming to establish a straightforward bias-aligned hypothesis. Then, it applies a visual explanation to the dataset using the amplified model. Finally, correlation-aware slicing is executed, a process in which clustering is conducted within the bias-amplified feature space. To enhance this process, we suggest integrating a visual explanation component $h$ before keyword generation.

### v3

@@ Abstract
The proliferation of machine learning models in critical decision-making processes has underscored the need for bias discovery and mitigation strategies. Identifying the reasons behind a biased system is not straightforward, since in many occasions they are associated with hidden spurious correlations which are not easy to spot. Standard approaches rely on bias audits performed by analyzing model performance in predefined subgroups of data samples, usually characterized by common attributes like gender or ethnicity when it comes to people, or other specific attributes defining semantically coherent groups of images. However, it is not always possible to know a priori the specific attributes defining the failure modes of visual recognition systems. Recent approaches propose to discover these groups by leveraging large vision language models, which enable the extraction of cross-modal embeddings and the generation of textual descriptions to characterize the subgroups where a certain model is underperforming. In this work, we argue that incorporating visual explanations (e.g. heatmaps generated via GradCAM or other approaches) can boost the performance of such bias discovery and mitigation frameworks. To this end, we introduce Visually Grounded Bias Discovery and Mitigation (ViG-Bias), a simple yet effective technique which can be integrated to a variety of existing frameworks to improve both discovery and mitigation performance. Our comprehensive evaluation shows that incorporating visual explanations enhances existing techniques like DOMINO, FACTS and Bias-to-Text, across several challenging datasets, including CelebA, Waterbirds, and NICO++.
@@ 1 Introduction
The increasing adoption of deep learning-based image classification models, notably in systems like facial recognition software, raises concerns related to biases and fairness[3]. Biases in such systems can lead to unfair and discriminatory outcomes, making it crucial to identify and address their underlying causes. Unfortunately, determining in which subgroups a given system may exhibit biased behavior is not an easy task. The typical approach to detect biases consists in auditing the system with respect to predefined attributes (often referred to as protected attributes), like gender, sex, age, or ethnicity, by comparing specific fairness metrics [2]. However, attributes like sex, age, or ethnicity are not always the real causes that explain the poor performance of a given model in a particular set of data samples. For example, in face recognition, a system may systematically fail to classify people with glasses due to the fact that they were not well represented in the training set. In this case, auditing with standard demographic attributes may not be able to reveal the real reason for bias. In other cases, such as the classification of animals or objects, the attributes that characterize the underperforming groups may not be obvious. Spurious correlations could lead to systems that make incorrect predictions because the model has learned these false relationships from the training data [8], or predictions which are actually correct but for the wrong reasons [24]. This issue often arises when the model is trained on datasets that do not accurately represent the real-world scenarios on which it is intended to work with, or when the dataset itself contains hidden biases. Thus, discovering the failure modes and providing explanations useful to understand the real reasons behind them becomes a crucial task when designing robust and fair systems. Existing literature refers to this task as unsupervised discovery of bias [14], or slice discovery [7], in the sense that such methods aim at mining the input data for semantically meaningful subgroups (slices) on which the model performs poorly. In this context, a slice is defined as a group of data samples that share a common attribute or characteristic that is not related to the target label.
@@
Recent studies have shown that text explanations [12] produced by multimodal large vision language models (VLMs) [6] can help in discovering and describing subpopulations on which a model systematically underperforms, without the need for protected attribute annotations. At the same time, a different line of work [14] has qualitatively demonstrated that visual explanations [1] derived from biased models, especially for instances originating from conflicting data slices, predominantly highlight spurious features usually associated with shortcut learning. In this work, we hypothesize that since visual explanations for biased models tend to underscore spurious correlations, they can serve as instrumental tools in informing VLMs when they are used to identify and mitigate undesired biases. To this end, we leverage the descriptive power of recent multimodal VLMs, and combine it with the capabilities of visual explanation mechanisms to uncover biases contributing to systematic failures for unknown slices in visual recognition models. Our extensive empirical evaluation highlights the advantages of the proposed method, being easily adapted to different bias discovery frameworks and systematically boosting their performance.
@@ 2 Related work
Visual explanations. Visual explanations in computer vision have largely contributed to the interpretability and trustworthiness of deep learning models. They provide insights into the decision-making process of neural networks by highlighting regions within an input image that are significant for predictions. One prominent technique in this domain is Gradient-weighted Class Activation Mapping (Grad-CAM) [21]. Grad-CAM utilizes the gradients flowing into the final convolutional layer of a convolutional neural network to understand which features are most important for a particular decision. By creating a heatmap of these weighted features, Grad-CAM offers a visual representation that demonstrates how the model focuses on different parts of the image to make a decision. This method not only aids in improving model transparency but also helps in diagnosing potential reasons for misclassifications. GradCAM has previously been employed as a description mechanism to interpret failure modes in [14]. However, even though the work of Krishnakumar et al [14] shows the potential of visual explanations for this task, they mostly provide qualitative examples and do not propose an automatic approach to bias discovery and mitigation. Singla and co-workers [22] present another interesting work where they generate annotations for visual explanations of the ImageNet dataset. They released a dataset called Salient ImageNet, which provides heatmaps distinguishing between spurious and core features, where the last ones correspond to a set of visual features that are always a part of the object definition. Even though they employ human supervision to identify the visual explanations related to spurious correlations (making the process difficult to automate), this work shows that such heatmaps can actually be used to pinpoint this type of spurious attributes. Here, we plan to leverage visual explanations (like GradCAM heatmaps and other variants) as a mechanism to direct the attention of bias discovery and mitigation methods towards spurious correlations.
@@
Automatic bias discovery using VLMs. Several recent efforts have been conducted in the identification of biased data slices using cross-modal embeddings. The DOMINO method [7] combines cross-modal embeddings and error-aware mixture models to find semantically coherent clusters of images where a given classification system is failing. Textual descriptions derived from the cross-modal embeddings are then used to characterize the failure modes that were just discovered. A similar approach is introduced by the FACTS (First Amplify and then Slice to Discover Bias) [27] method, where the authors propose to first amplify the sources of bias in the training to ease its identification. This process is carried out in two stages. The first stage involves amplifying the bias by regularizing the model. This is achieved through regularization by specifically increasing the weight decay factor $\lambda$, which penalizes the most significant weights in the model. Regularization is a common technique used to prevent overfitting, but in this context, it serves a dual purpose: it not only helps to prevent overfitting but also forces the model to rely more on the spurious attribute by simplifying the hypothesis space. The second stage consists of identifying underperforming data segments that exhibit unique correlations by employing mixture modeling within a feature space aligned with bias, a method known as correlation-aware slicing, similar to the aforementioned DOMINO. In [12], a different approach called Bias-to-Text (B2T) was introduced. B2T uses linguistic interpretation to identify and mitigate biases in vision models, including image classifiers. By generating linguistic descriptions of images using VLMs, B2T extracts keywords indicative of bias from the wrongly predicted examples, enabling a more focused understanding and correction of prejudices within the models.
@@
Bias mitigation: improving robustness to spurious correlations. Discovering conflicting slices is important, but mitigating such biases is also crucial. A common approach is based on group distributionally robust optimization (GroupDRO) [20]. GroupDRO aims to enhance model performance across different groups or subpopulations in the data, based on predefined protected attributes like race or gender. The objective is to minimize the worst-case loss across these groups, thereby encouraging fairness and reducing disparities in model performance among them. However, one notable drawback of this technique is the requirement to assign specific group labels to every dataset entry, which can lead to substantial increases in annotation costs, and requires to know before-hand such groups. Another existing approach in the bias mitigation field is JTT (Just Train Twice) [15]. JTT starts with the standard training of the model to pinpoint examples that have been misclassified. Following this, the training process is adjusted by reweighting the dataset, giving greater importance to these previously misclassified examples. JTT forces the model to pay more attention to examples where the model is failing, thereby enhancing its performance on those specific examples. Here, we will employ JTT as a baseline approach for comparison, and propose ways to integrate visual explanations into the group definitions for GroupDRO.
@@ 3 Preliminaries and Problem Formulation
The dataset $\mathcal{D}$ is partitioned into different and not equally distributed groups or slices. A group is defined based on the combination of the target labels $y\in\mathcal{Y}$ and the spurious attributes $a_{i}\in\mathcal{A}$ that spuriously correlates with the label (i.e. $\mathcal{G}=\mathcal{A}\times\mathcal{Y}$). Formally, given a pair $(a_{i},y^{\star})$, we define a group $g(a_{i},y^{\star})$ as the set of data samples labeled with $y^{\star}$ which contain attribute $a_{i}$, i.e. where $a_{i}(x)=1$.
For example, in the CelebA dataset[17], where the task involves classifying people with blond hair, the spurious attribute is the gender, as there is a higher frequency of blond women in the dataset. Figure 1 illustrates this issue, and a similar case found in the Waterbirds dataset [20], highlighting the classes with the least training samples per group.
@@ 4 ViG-Bias: Visually Grounded Bias Discovery and Mitigation
Consider a dataset $D$ with samples $(x,y,a)\in\mathcal{D}$ drawn from $\mathcal{X}\times\mathcal{Y}\times\mathcal{A}$, where $a$ is spuriously correlated with the label $y$. We are interested in discovering the visual features upon which a model has relied to make a prediction. Prior work [11, 22] suggests that, under the presence of spurious correlations, visual explanations for models trained with standard empirical risk minimization (ERM) tend to focus on such shortcuts, while ignoring the core features. We thus resort to GradCAM heatmaps generated using an image classifier that was trained with standard ERM on the dataset of interest $\mathcal{D}$. Such heatmaps will then be used as visual explanations to guide the attention of bias discovery methods (alternative visual explanation methods like ScoreCAM [26], FullGrad [23], GradCAM++ [4] are also evaluated in Section 6.3). GradCAM heatmaps are obtained by first computing the gradients of the target class with respect to the feature maps of the last convolutional layer, as neurons in this layer offer best compromise between high-level semantics and detailed spatial information [21]. These gradients flowing back are then globally-averaged over the width and height dimensions to obtain the importance weights of every feature map in this last layer. Then, the ReLU activation function is applied on the weighted average of the feature maps, so that only the features that have a positive influence on the class of interest are finally highlighted. These heatmaps are finally normalized between 0 and 1.
@@ 4.1 Are Visual Explanations Focusing on Spurious Correlations?
The work of Krishnakumar [14] and colleagues provides qualitative examples showing that GradCAM heatmaps for biased models tend to focus on spurious correlations. Here we perform a more systematic quantitative evaluation to confirm this hyphotesis. To this end, we design a simple experiment using CelebA (where the conflicting slices are well known) that consists of three steps:
@@
    We first construct binary segmentation masks $m^{c}_{i}$ and $m^{s}_{i}$ for all images $x_{i}$ in both datasets, segmenting the core features associated with the object of interests ($m^{c}_{i}$ representing, for example, the hair in CelebA), and the spurious attribute ($m^{s}_{i}$ representing the face) as shown in Figure 1. We do so by using LangSAM  [^1] https://github.com/luca-medeiros/lang-segment-anything, an open-source model that combines Segment Anything Model (SAM) [13] with GroundingDINO [16] to enable instance segmentation from text prompts.
@@
Figure 2: Visual explanation methods (e.g. GradCAM[21] for the spurious attribute blond) help with identifying spurious correlations. We use the Intersection-over-Union (IoU) metric to measure the percentage of overlap between a binary image representing the spurious feature and the image we get after applying a visual explanation method.
@@ 4.2 Improving Unsupervised Bias Discovery via Visual Explanations
We modify DOMINO by simply pre-processing the images before the cross-modal embeddings are obtained, i.e. we generate heatmaps using the pretrained classifier and extract embeddings for $h(x)$ instead of $x$. Note that other types of visual explanation methods could also be used, as discussed in Section 6.3.
Visually Grounded FACTS (ViG-FACTS): The FACTS method is similar to DOMINO, but it consists of a two-stages process, where biases are first amplified and then slicing is performed. In the first stage, we increase the model’s reliance on spurious, often bias-indicative correlations, which are amplified using regularization. In general, regularization helps in preventing overfitting and encourages the model to learn more robust patterns that are assumed to be applicable beyond the training set. However, in FACTS, the regularization is applied to constrain the model’s capacity and learn the bias-aligned slices where spurious correlation holds. This leads to the model developing a strong dependency on these correlations, making the biases more pronounced and identifiable. Once these biases are amplified and more clearly delineated, we proceed to the second stage -slicing- where an error-aware mixture model that uses the CLIP embeddings similarly to DOMINO is applied.
@@
Figure 4: Our objective is to pinpoint slices of data where a spurious correlation exists between a task-irrelevant attribute (such as gender) and the label of interest. For instance, in the given example, women who are not blond correspond to a bias-conflicting slice, while not blond men form a bias-aligned slice. FACTS amplifies correlations, aiming to establish a straightforward bias-aligned hypothesis. Then, it applies a visual explanation to the dataset using the amplified model. Finally, correlation-aware slicing is executed, a process in which clustering is conducted within the bias-amplified feature space. To enhance this process, we suggest integrating a visual explanation component $h$ before keyword generation.